OpenStudy (anonymous):

Medal If y = 5x +8 , write x= f (y)

OpenStudy (anonymous):

y = 5x +8 Subtract -5x from both sides -5x + y = 8 Subtract y from both sides -5x = -y +8 Divide both sides by -5 That should give you your answer. :)

OpenStudy (anonymous):

how would you solve this

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

i got -1.6 as my answer am i right

OpenStudy (anonymous):

Tell me how you got 1.6

OpenStudy (anonymous):

-1.6*

OpenStudy (anonymous):

8/-5 = 1.6

OpenStudy (anonymous):

Do you only need the y value?

OpenStudy (anonymous):

i need the x value

OpenStudy (anonymous):

So you are isolating x, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You just divide the rest of the equation by -5, and x is then isolated. x = -1/5y - 8/5

OpenStudy (anonymous):

Unless your teacher wants you to, you don't have to solve the fractions either.

OpenStudy (anonymous):

-1/5y = -0.2 -1.6

OpenStudy (anonymous):

Where did you get the -0.2 and -1.6?

OpenStudy (anonymous):

I know how you got -1.6

OpenStudy (anonymous):

-1 divded by 5 = -0.2

OpenStudy (anonymous):

You have to keep your variables in the equation, so you would have... x = -0.2y - 1.6

OpenStudy (anonymous):

Unless you are given another equation or a set of coordinates, you can't do much more to the equation given, so x = -1/5y - 8/5 or x = -0.2y - 1.6 would be your final answer.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You're very welcome. :)